OpenStudy (anonymous):

Find the domain, intercepts, the intervals where function is increasing and decreasing, extreme values, intervals of concavity, points of inflection, asymptotes (vertical, horizontal, or slant). Show the1st and the 2nd derivative tables. Sketch a graph. f(x)= x^2+12/(x-1)

OpenStudy (zzr0ck3r):

lol thats alot, which one are you having problems with?

OpenStudy (anonymous):

all of it :)

OpenStudy (zzr0ck3r):

if you dont know how to find the domain of a function I would suggest not taking a calc class

OpenStudy (zzr0ck3r):

to find the domain we want to find every number that x CAN be so we look for what x CANT be for the most part...you cant devide by zero or take the sqrt of a negative number so when asked a question about domain look for the oppertunity for both of those things here we can devide by zero because we have a ration with a variable in the bottom 12/(x-1) so the only "problem" here is that x-1 so what values of x make this bad? well set the bottom = to zero and solve x-1=0 iif x = 1 so your domain is all real numbers except 1

OpenStudy (zzr0ck3r):

when a function is increasing its derivative is positive and when decreasing its derivative is negative so first off we need the derivative... to take the derivative we take the derivative of each term and add them to each other so d(x^2)/dx + d(12/(x-1))/dx the second part is the quotent rule for derivatives or the chain rule..how ever you want to do it. d(x^2)/dx + d(12/(x-1))/dx = 2x + d(12(x-1)^-1)dx = 2x - 12(x-1)^(-2) = 2x - 12/(x-1)^2

OpenStudy (zzr0ck3r):

does this make since?

OpenStudy (zzr0ck3r):

i guess get back to me when you have time:P